Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability: Amy and her deck of cards

Expert replies
by tabsang » Mon Dec 24, 2012 2:17 am
Amy has two decks of 52 cards each: Deck 1 and Deck 2. She takes 8 black cards from Deck 2 and adds them to Deck 1 and shuffles it thoroughly. She now picks a card from the newly formed pack of cards. If the probability of either picking a red ace or a king from the newly formed pack is greater than 1/8, what is the probability that Amy picks a black king or a red Jack from the new pack?

(A) 1/6
(B) 1/8
(C) 1/9
(D) 1/10
(E) 1/12

OA: [spoiler](A)[/spoiler]
But I got a different answer.
Join the discussion
Source: — Problem Solving |

by tabsang » Mon Dec 24, 2012 2:30 am
Here's my approach:

Let "k" be the number of kings (black) that may be part of the 8 BLACK cards added to Deck 1.
No. of red aces in the new deck remains unchanged i.e. 2
No. of cards in the new deck = 52+8 = 60
No. of kings in the new deck = 4+k (2 Red kings + 2 Black kings + k Black kings)

Thus, probability of picking a red ace or a king is :

2/(52+8) + (4+k)/(52+8) = (6+k)/60

Now, since it's given that the above probability is greater than 1/8
We get,
(6+k)/60 > 1/8

i.e. k > 3/2
i.e. k > 1

k can only take 3 values (0 black kings or 1 black king or 2 black kings)
Thus, since k>1, k=2.

Now, the required probability of Amy picking a black king or a red jack from the new pack can be calculated as follows:

Total no. of black kings in the new pack = 4 (2 from Deck 1 + 2 from the 8 cards added to Deck 1)
Total no. of red Jacks in the new pack = 2

Thus,

4/60 + 2/60 = 6/60 = 1/10.

I got (D) as the answer.
Did anyone else get the same answer???

Cheers,
Taz
Join the discussion

by The Iceman » Mon Dec 24, 2012 2:47 am
tabsang wrote:I got (D) as the answer.
Did anyone else get the same answer???

Cheers,
Taz
D seems to be the right ans. I have a two liner solution for this:

If there are k kings in the pack, then 2+k /60 > 1/8 => k>=6 => k=6 and both the black kings have been borrowed from deck 2.

Required probabilty = (4+2)/60 = 1/10
Join the discussion

by anirudhadhopate » Wed Dec 26, 2012 9:13 am
IMO D should be the answer. What is the answer?

Btw what is source of this problem?
Join the discussion

by tabsang » Wed Dec 26, 2012 9:34 am
anirudhadhopate wrote:IMO D should be the answer. What is the answer?

Btw what is source of this problem?
This from a printed set of questions that I have.
The OA is (A) but I've verified that the correct answer is infact (D).

So Cheers, You got it spot on :),
Taz
Join the discussion